OpenStudy (anonymous):

Integral x*ln(x+1) - x*ln(x) dx

OpenStudy (luigi0210):

Is it this? \[\LARGE \int(x*ln(x+1)-x*lnx)~dx\]

OpenStudy (anonymous):

exactly (;

ganeshie8 (ganeshie8):

by parts ?

OpenStudy (anonymous):

Yeah looks like it.....and I think the second one can be looked up on table of integrals. Ill keep playing with it.

ganeshie8 (ganeshie8):

simplify a bit before doing by parts

ganeshie8 (ganeshie8):

\(\LARGE \int x*\ln(x+1)-x*\ln x~dx\) \(\LARGE \int x[\ln(x+1)-\ln x]~dx\) \(\LARGE \int x\ln(1+\frac{1}{x})~dx\)

ganeshie8 (ganeshie8):

u = x^2/2 v = ln(1+1/x)

ganeshie8 (ganeshie8):

should be easy

OpenStudy (anonymous):

Thanks!!!!!

ganeshie8 (ganeshie8):

np :)